OpenStudy (anonymous):

Ok i need some help with this problem 1/1+x^n + 1/1+x^-n Simplify your answer I can't seem to get the correct answer

OpenStudy (anonymous):

\[\huge\frac{1}{1+x^n}+\frac{1}{1+x^{-n}}\]??

OpenStudy (aum):

Is the problem: \[ \frac{1}{1+x^n} + \frac{1}{1 + x^{-n}}~~ ? \]

OpenStudy (anonymous):

yes

OpenStudy (aum):

Multiply the top and bottom of the second one by x^n and then add the two.

OpenStudy (anonymous):

ok

OpenStudy (aum):

\[\large \frac{1}{1+x^n} + \frac{1}{1 + x^{-n}} = \frac{1}{1+x^n} + \frac{1}{1 + x^{-n}} * \frac{x^n}{x^n} = \\ \large \frac{1}{1+x^n} + \frac{x^n}{x^n + 1} = \frac{1+x^n}{1+x^n} = 1 \]

OpenStudy (anonymous):

a miracle will occur

OpenStudy (anonymous):

ok that makes so much more sense now thank you

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

can you help with another problem as well or are you busy?

OpenStudy (aum):

go ahead.

OpenStudy (anonymous):

ok let me copy it real quick

OpenStudy (anonymous):

can you open that

OpenStudy (aum):

No link to click on; nothing to open.

OpenStudy (anonymous):

ok let me add another one

OpenStudy (anonymous):

try that

OpenStudy (aum):

Factor 5x out of the numerator. Try to factor the quadratic that you get. Try to factor the quadratic in the denominator. Tell me what you get.

OpenStudy (anonymous):

ok ill try that

OpenStudy (anonymous):

ok so I'm left with 5x(x^2-x-2) on the top and the bottom I'm getting x+2 4x+2 which would give me back the equation but with a positive 13 and thats not right

OpenStudy (aum):

The (x^2-x-2) in the numerator can be factored further. The denominator has not been factored properly.

OpenStudy (anonymous):

ok yes i see that it can go further ill try it again

OpenStudy (anonymous):

the numerator would be 5x(x+1)5x(x-2) and the denominator would be the same thing i got but with negative signs i see what i did wrong

OpenStudy (aum):

numerator is correct.

OpenStudy (anonymous):

the denominator would be 4x-5 times x-2

OpenStudy (aum):

correct. And the (x-2) will cancel with the same in the numerator.

OpenStudy (anonymous):

yes i see that

OpenStudy (anonymous):

so I'm left with 5x(x+1) on the top and 4x-5 on the bottom correct?

OpenStudy (aum):

Yes. \[ \large \frac{5x(x+1)}{(4x-5)} \]

OpenStudy (anonymous):

ok thank you for your help

OpenStudy (aum):

you are welcome.